A scientist wishes to compress 50.0 ml of ammonia gas at normal atmospheric pressure to a volume of 1.0 ml. What pressure in atmospheres is necessary
1
Expert's answer
2012-12-12T08:24:15-0500
Using Law: PV = nRT we can calculate P (pressure)

But this law becomes P1V1/T1 = P2V2/T2 for one compound under different conditions

T1 = T2 = const, so P1V1 = P2V2

P1 = 1 atm (STP), so P2 = P1V1/V2 = 1*50.0/1.0 = 50 atm